Find mass of sodium chloride (in Ib) required to prepare 100 gallons of 2.0 M solution of sodium chloride
1
Expert's answer
2012-10-10T10:34:13-0400
2.0 M means 2 mol per liter

3.785 liters in 1 gallons

378.5 liters
in 100 gallons


2 mol -------------- 1 L
x
-----------------378.5 L

x=757 mol

n =m/Mm

m= n * Mm= 757 *
58.5 = 44284.5 g =44.2845 kg